# A note on the cohomology of moduli spaces of local shtukas

### v1

@@ 4.2. Vanishing
One interesting aspect of Theorem 3.5 is its uniformity when varying $\mu_{\bullet}$: Only the $W_{E_{\bullet}}$-representations change [^4] A trivial consequence is that, for fixed $(G,b)$ and $\rho$, there exists an $N$ such that $H^{i}(R\Gamma(G,b,\mu_{\bullet})[\rho])$ vanishes for all $|i|>N$ and all $\mu_{\bullet}$.. This observation can sometimes be used to propagate vanishing results for some $\mu_{\bullet}$ to a larger collection (Conjecture 1.1(2) was recently proven by the first author in [Han21] in many cases where $\mu_{\bullet}$ is a single minusucle cocharacter). To illustrate the method, we reprove the strong Kottwitz conjecture for inner forms of $\operatorname{GL}_{n}$, which was previously proved by the first author [Han21, Thm. 1.9], using results on averaging functors from [ALB21].
@@
Assume that $G$ is an inner form of $\operatorname{GL}_{n}$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\pi\boxtimes V_{\varphi}$, where the right hand is concentrated in degree 00 and $\pi$ is the irreducible admissible representation of $G(E)$ corresponding to $\varphi$ under the usual local Langlands correspondence.
@@
Note that we are in the stable situation here, so we need to show that $i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)$ is $\pi$ concentrated in degree 00. That it is a shift of $\pi$ follows from [HKW22, Thm. 1.0.3], since its Fargues–Scholze $L$-parameter is $\varphi$. To prove vanishing, note that we can choose a minuscule cocharacter $\mu_{b}$ such that $R\Gamma(G,b,\mu_{b})[\rho]=i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)\boxtimes W_{\varphi}$, where $W$ the dual of the irreducible representation with extreme weight $\mu_{b}$. By [Han21, Thm. 1.6], $R\Gamma(G,b,\mu_{b})[\rho]$ vanishes outside degree 00, finishing the proof. ∎
@@
as objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ (i.e. forgetting the $W_{E_{\bullet}}$-action), where we have used that $i_{1}^{\ast}=i_{1}^{!}$, $i_{b!}\rho^{\vee}=i_{b\ast}\rho^{\vee}$, and the interplay between $\mathbb{D}$ and pullback/pushforward, and the Hecke operators [FS21, Thm. IX.0.1(i)]. This gives us the following vanishing theorem. In its formulation, we note that there is a canonical isomorphism between the cocenters of $G$ and $G_{b}$ over $E$, so any smooth character of $G(E)$ can be naturally viewed as a smooth character of $G_{b}(E)$ (and vice versa).
@@
Assume that the $\delta_{i}$ are one-dimensional and that, writing $i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}[n_{i}]$, the $\pi_{i}$ are distinct. Assume further that there is a smooth character $\chi$ of $G(E)$ such that $\rho^{\vee}\cong\rho\otimes\chi$ and $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ for all $i$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is concentrated in degree 00.
@@
Since $\rho^{\vee}\cong\rho\otimes\chi$, Verdier duality gives us that $\mathbb{D}(R\Gamma(G,b,\mu_{\bullet})[\rho])\cong R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]$. Computing from the definitions, one sees that $R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]\cong R\Gamma(G,b,\mu_{\bullet})[\rho]\otimes\chi$. Thus, evaluating both sides using Theorem 3.8, we get that
@@
Since $\pi^{\vee}_{i}\cong\pi_{i}\otimes\chi$ and the $\pi_{i}$ are assumed to be distinct, it follows that $n_{i}=0$ for all $i$ as desired. ∎
@@
    In the stable case, the assumption $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ can be dropped, and instead deduced as a consequence of the argument.
@@
    That the $\pi_{i}$ are distinct follows from the refined local Langlands correspondence whenever $G$ is an extended pure inner form of its quasi-split inner form. It may fail when this is not the case, for example when $G$ is the non-split inner form of $\operatorname{SL}_{2}$ and $S_{\varphi}=(\mathbb{Z}/2)^{2}$, in which case the $\pi_{i}$ are all equal.

### v2

@@ 4.2. Vanishing
Recall that we assume Setup 4.1 throughout §4. One interesting aspect of Theorem 3.8 is its uniformity when varying $\mu_{\bullet}$: Only the $W_{E_{\bullet}}$-representations change [^7] A trivial consequence is that, for fixed $(G,b)$ and $\rho$, there exists an $N$ such that $H^{i}(R\Gamma(G,b,\mu_{\bullet})[\rho])$ vanishes for all $|i|>N$ and all $\mu_{\bullet}$.. This observation can sometimes be used to propagate vanishing results for some $\mu_{\bullet}$ to a larger collection [^8] Conjecture 1.1(2) was recently proven by the first author in [Han21] in many cases where $\mu_{\bullet}$ is a single minusucle cocharacter.. To illustrate the method, we reprove the strong Kottwitz conjecture for inner forms of $\operatorname{GL}_{n}$, which was previously proved by the first author [Han21, Thm. 1.9], using results on averaging functors from [ALB21].
@@
Assume Setup 4.1 and that $G$ is an inner form of $\operatorname{GL}_{n}$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]\cong\pi\boxtimes V_{\varphi}$, where the right hand is concentrated in degree 0 and $\pi$ is the irreducible admissible representation of $G(E)$ corresponding to $\varphi$ under the usual local Langlands correspondence.
@@
Note that we are in the stable situation here, so we need to show that $i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)$ is $\pi$ concentrated in degree 0. That it is a shift of $\pi$ follows from [HKW22, Thm. 1.0.3], since its Fargues–Scholze $L$-parameter is $\varphi$. To prove vanishing, note that we can choose a minuscule cocharacter $\mu_{b}$ such that $R\Gamma(G,b,\mu_{b})[\rho]=i^{1\ast}\operatorname{Act}_{\chi_{b}}(i^{b}_{\ast}\rho)\boxtimes W_{\varphi}$, where $W$ the dual of the irreducible representation with extreme weight $\mu_{b}$. By [Han21, Thm. 1.6], $R\Gamma(G,b,\mu_{b})[\rho]$ vanishes outside degree 0, finishing the proof. ∎
@@
as objects in $\mathcal{D}_{\mathrm{lis}}(\mathrm{Bun}_{G})$ (i.e. forgetting the $W_{E_{\bullet}}$-action), where we have used that $i_{1}^{\ast}=i_{1}^{!}$, $i_{b!}\rho^{\vee}=i_{b\ast}\rho^{\vee}$ (by [Han21, Thm 1.3] and [FS21, Thm IX.05(iv)], since $\varphi$ is elliptic), and the interplay between $\mathbb{D}$ and pullback/pushforward, and the Hecke operators [FS21, Thm. IX.0.1(i)]. This gives us the following vanishing theorem. In its formulation, we note that there is a canonical isomorphism between the cocenters of $G$ and $G_{b}$ over $E$, so any smooth character of $G(E)$ can be naturally viewed as a smooth character of $G_{b}(E)$ (and vice versa).
@@
Assume Setup 4.1, that the $\delta_{i}$ are one-dimensional, and that, writing $i^{1\ast}\operatorname{Act}_{\delta_{i}}(i^{b}_{\ast}\rho)=\pi_{i}[n_{i}]$, the $\pi_{i}$ are distinct. Assume further that there is a smooth character $\chi$ of $G(E)$ such that $\rho^{\vee}\cong\rho\otimes\chi$ and $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ for all $i$. Then $R\Gamma(G,b,\mu_{\bullet})[\rho]$ is concentrated in degree 0.
@@
Since $\rho^{\vee}\cong\rho\otimes\chi$, Verdier duality gives us that $\mathbb{D}(R\Gamma(G,b,\mu_{\bullet})[\rho])\cong R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]$. Computing from the definitions, one sees that $R\Gamma(G,b,\mu_{\bullet})[\rho\otimes\chi]\cong R\Gamma(G,b,\mu_{\bullet})[\rho]\otimes\chi$. Thus, evaluating both sides as a $G(E)$-representation using Theorem 3.8 and Corollary 3.11, we get that
@@
Since $\pi^{\vee}_{i}\cong\pi_{i}\otimes\chi$ and the $\pi_{i}$ are assumed to be distinct, it follows that $n_{i}=0$ for all $i$ with non-zero multiplicity, as desired. ∎
@@
    If $\varphi$ is stable, the assumption $\pi_{i}^{\vee}\cong\pi_{i}\otimes\chi$ can be dropped, and instead deduced as a consequence of the argument.
@@
    If, in addition to the assumptions in Proposition 4.8, we also assume Setup 4.3, then the assertion that the $\pi_{i}$ are distinct is equivalent to $\Pi_{\phi}(G)$ having size $r$. For example, this is true if $S_{\phi}/Z(\widehat{G})^{Q}\cong S_{\varphi}/Z(\widehat{G})^{Q}$ and $G$ is quasisplit (then $\Pi_{\phi}(G)$ is in bijection with $\operatorname{Irr}(S_{\varphi}/Z(\widehat{G})^{Q})$, which is equal to $\{\delta_{i}\delta_{j}^{-1}\mid i=1,\dots,r\}$ for arbitrary $j$). On the other hand, if $G$ is the non-split inner form of $\operatorname{SL}_{2}$ and $S_{\varphi}=(\mathbb{Z}/2)^{2}$, then we know that $\phi=\varphi$ and that $\Pi_{\phi}(G)$ has size $1$.